OpenStudy (anonymous):

multiply the polynomial below:

OpenStudy (anonymous):

\[\frac{ x^2-2x-8 }{ 6x^2-24x } \times \frac{ 2x^2-3x-2}{ x^2-4 }\]

OpenStudy (anonymous):

i imagine your job is to factor as much as you can and then cancel

OpenStudy (anonymous):

ok hold on let me try

OpenStudy (anonymous):

for example \(x^2-2x-8=(x-4)(x+2)\) and you will also get a factor of \((x+2)\) in the denominator because \(x^2-4=(x+2)(x-2)\)

OpenStudy (anonymous):

\[\frac{ 2x^2-7x-4 }{ x^2-4}\]

OpenStudy (anonymous):

i think you can do more

OpenStudy (anonymous):

-7x?

OpenStudy (anonymous):

-7x-1?

OpenStudy (anonymous):

ok lets go ahead and factor

OpenStudy (anonymous):

wait 2-7x +1 3-7x

OpenStudy (anonymous):

\[\frac{ x^2-2x-8 }{ 6x^2-24x } \times \frac{ 2x^2-3x-2}{ x^2-4 }\] \[\frac{(x-4)(x+2)(x-2)(2x+1)}{6x(x-4)(x+2)(x-2)}\]

OpenStudy (anonymous):

cancel away and you are not left with much just \[\frac{2x+1}{6x}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw